# DiffHO: Diffusion-Based Synthesis of Hand-Object Interactions from Textual Descriptions

### v1

@@ Abstract.
Generating natural hand-object interactions in 3D is challenging as the resulting hand and object motions are expected to be physically plausible and semantically meaningful. Furthermore, generalization to unseen objects is hindered by the limited scale of available hand-object interaction datasets. We propose DiffHO, a novel method to synthesize realistic, one or two-handed object interactions from provided text prompts and geometry of the object. The method introduces three techniques that enable effective learning from limited data. First, we decompose the task into a grasping stage and a text-based interaction stage and use separate diffusion models for each. In the grasping stage, the model only generates hand motions, whereas in the interaction phase both hand and object poses are synthesized. Second, we propose a compact representation that tightly couples hand and object poses. Third, we propose two different guidance schemes to allow more control of the generated motions: grasp guidance and detailed textual guidance. Grasp guidance takes a single target grasping pose and guides the diffusion model to reach this grasp at the end of the grasping stage, which provides control over the grasping pose. Given a grasping motion from this stage, multiple different actions can be prompted in the interaction phase. For textual guidance, we contribute comprehensive text descriptions to the GRAB dataset and show that they enable our method to have more fine-grained control over hand-object interactions. Our quantitative and qualitative evaluation demonstrates that the proposed method outperforms baseline methods and leads to natural hand-object motions. Moreover, we demonstrate the practicality of our framework by utilizing a hand pose estimate from an off-the-shelf pose estimator for guidance, and then sampling multiple different actions in the interaction stage. The project page along with videos can be found at https://diffh2o.github.io/
@@ 1. Introduction
Modeling and understanding hand-object interactions is an important and crucial task to empower machines to interact with and assist humans. Being able to seamlessly generate hand-object motions holds the promise to enable synthetic data generation (Leonardi et al., 2023), assist robots in training human-robot interactions in simulation, or enhance realism in virtual manipulation experiences. For example, in virtual reality (VR), hand interactions still often rely on heuristics and controllers that simply attach objects to the hand according to pre-defined grasps. Being able to faithfully produce object manipulation based on an input signal such as text or a few past frames of hand and object poses could largely increase the immersiveness of such interactions.
@@
Generating realistic hand-object interactions in 3D is challenging as the resulting motions are required to be plausible in several different aspects. First, the motions must be plausible in terms of geometry, such hand and object intersections are minimized and the grasp appears stable. Second, the motions must be plausible in terms of semantics, such as the hands respect natural object affordances (e.g., a cup is grasped by its handle and not flipped upside down). Third, the motions must be plausible in terms of time, such as the hand and object motions are synchronized and the dynamics appear natural. Another challenge in generating hand-object interactions comes from the limited scale of existing hand-object datasets, which are around 10x smaller than human motion datasets (Mahmood et al., 2019) and 1000x smaller than image datasets (Deng et al., 2009).
@@
Several recent works have successfully leveraged diffusion models for full-body human motion generation but either neglect objects (Tevet et al., 2023; Karunratanakul et al., 2023) or focus only on coarse motion of larger objects (e.g., a chair) without finger predictions (Xu et al., 2023). Another method, IMoS (Ghosh et al., 2023), can generate hand-object interactions with smaller objects (e.g., a cup or stapler) by first optimizing for upper body motions and then optimizing for object poses in a post-processing step. However, IMoS assumes that the object is in hand from the beginning, can handle only objects seen during training, and, as shown in our qualitative evaluation, the optimization often yields artifacts (e.g., hand-object interpenetration or non-plausible contact points). In this paper, we aim to learn a model that can generate natural, fine-grained hand-object interactions and generalize to objects unseen during training.
@@
Towards this goal, we contribute a novel diffusion-based framework, DiffHO, that can generate hand-object interactions based on provided text prompts and geometry of the object. We propose several techniques to deal with the challenges of data scarcity and object generalization. Decoupling Grasping and Interaction: Hand-object interactions can be invariably split into a grasping phase where the hand(s) approaches the static object for grasping, and an interaction phase where the object is manipulated based on an action intent. For example, drinking from a cup can be performed with multiple grasps, and given a grasp, multiple actions are possible. We use this simple observation to obtain a grasping and an interaction diffusion model, and introduce an inpainting technique called subsequence guidance to allow continuity between the two outputs. Different from previous diffusion frameworks that employ coarse-to-fine (2D trajectory-to-full motion) two-stage generation (Karunratanakul et al., 2023), we split the task temporally. Pose Representation: We introduce a compact representation that models hands in parametric space. To couple the hands to the object, we represent the global hand positions relative to the normalized object pose at the initial frame and further include distance information between the 3D hand joints and the object surface. Unlike previous representations utilized in human motion diffusion (Tevet et al., 2023) and hand-object interaction (Christen et al., 2022), we find a non-redundant representation directly in parametric space to perform best. Controllability: We provide two ways to add more control to our diffusion models’ output. First, we propose control through fine-grained textual descriptions. To this end, we contribute textual descriptions to the GRAB dataset and show that they enable more robustness to unseen test prompts and increase controllability, e.g., by allowing to define the interacting hand and the action to perform through the input prompts. Second, in grasp guidance, we leverage a single grasp reference which connects the grasping and interaction phase and use it to guide our two diffusion models at inference time. This grasp pose can either be obtained from motion capture or from an image-based pose estimator (Pavlakos et al., 2023) as we show in our experiments Section 5.8.
@@
To demonstrate the effectiveness of our approach, we run perceptual studies and experiments on the GRAB dataset (Taheri et al., 2020). We first compare against IMoS (Ghosh et al., 2023). Our results indicate that our method generates better results across a variety of physics-based and motion metrics. To justify our technical contributions, we compare our final model against human-body diffusion baselines by adapting these methods to hand-object interactions. We show that we generate motions of higher quality while enabling more control of the outputs. Furthermore, we demonstrate that detailed textual descriptions increase robustness to unseen texts and offer better control of the diffusion model than training with heuristics-generated text descriptions. Decoupling and grasp-guidance brings high practical value to DiffHO. We show this with the first application of using an image-based pose/shape estimator (Pavlakos et al., 2023) to guide the diffusion to the grasp, and thereafter sampling multiple different actions through text. We prove the utility of our method by largely outperforming the state-of-the-art on a standard benchmark and in perceptual studies.
Our contributions can be summarized as follows:
To the best of our knowledge, DiffH2O is the first method that synthesizes hand-object interactions on unseen objects from textual descriptions, unlike previous generative models for human-body motion and seen-object HOI synthesis.
We propose a two-stage diffusion process that splits generation into grasping and action-based interaction.
To increase the controllability of outputs and improve generalization to unseen objects, we propose grasp guidance and subsequence guidance that can be applied at inference time to the diffusion process.
We contribute detailed textual annotations to the GRAB dataset along with experiments showing that these textual descriptions increase robustness and enable fine-grained control of the model output.
Code and data will be made public upon acceptance.
@@ 2.1. Hand-Object Synthesis
Recent efforts in interaction synthesis have largely been driven by the surge of high-quality human-object interaction datasets (Taheri et al., 2020; Brahmbhatt et al., 2019; Hampali et al., 2020; Brahmbhatt et al., 2020; Chao et al., 2021; Kwon et al., 2021; Fan et al., 2023; Liu et al., 2022). Some studies focus on generating coarse full-body object interactions (Zhang et al., 2022a; Wang et al., 2021; Lee and Joo, 2023; Hassan et al., 2023; Luo et al., 2021), such as carrying or moving around boxes. FLEX (Tendulkar et al., 2023) trains a hand and body pose prior and later optimizes the priors to achieve diverse, static full-body grasps. GOAL (Taheri et al., 2022) and SAGA (Wu et al., 2022) use CVAEs to generate approaching motions for full-body grasps, whereas TOHO (Li et al., 2024) models both approaching and manipulation tasks via neural implicit representations. Similarly, Braun et al. (2024) model the full range of motion, leveraging physics simulation and reinforcement learning. In contrast to these works, we focus on fine-grained hand-object interactions and generalization to unseen objects. Furthermore, we leverage language prompts as input to generate interaction sequences. Among methods that model full-body object interaction, closest to our work is IMoS (Ghosh et al., 2023), a two-stage method to generate hand-object interactions on seen objects based on language commands. Starting from a grasping state, they first generate body motions and then optimize for object trajectories using a heuristics-based optimization to model hand-object interactions. By contrast, our model directly predicts the hand and object poses, models both approaching and manipulation phases, and generalizes to unseen objects.
@@
Another line of research focuses on generation of hand-object interaction sequences in isolation from full-body due to wide applications in VR and the need for a dedicated model for fine-grained details of finger motion. One prominent solution is to turn to physical simulation and reinforcement learning (Christen et al., 2022; Mandikal and Grauman, 2021; Garcia-Hernando et al., 2020; Rajeswaran et al., 2018; Qin et al., 2021). A number of studies propose to learn dexterous manipulation tasks from full human demonstration data either collected via teleoperation (Rajeswaran et al., 2018) or from videos (Garcia-Hernando et al., 2020; Qin et al., 2021). Mandikal and Grauman (2021) propose a reward function that incentivizes dexterous robotic hand policies to grasp in the affordance region of objects. She et al. (2022) propose a new dynamic state representation to generate dexterous grasps. D-Grasp (Christen et al., 2022) leverages reinforcement learning and physics simulation to generate diverse hand-object interactions from sparse reference inputs. ArtiGrasp (Zhang et al., 2024) extends this to two-handed grasps and generates articulated object motions.
@@
To model hand-object interaction sequences, another solution is to rely on purely data-driven frameworks. For example, (Zhou et al., 2022; Anonymous, 2023) propose methods that enable denoising hand-poses from noisy sequences of hand-object poses. Ye and Liu (2012) predict the local hand pose given full body and object motion. Similarly, ManipNet (Zhang et al., 2021) predicts local hand poses for two-handed interactions based on wrist-object trajectories. Given object motion of articulated objects, CAMS (Zheng et al., 2023) predicts one-handed poses that align with the object motion, whereas Chen et al. (2023b) focus on grasping objects using dexterity in the environment. All these works either assume a hand-object trajectory, focus only on one hand, ignore semantic action information, or do not generalize well to unseen objects. In contrast, we propose a framework that allows the synthesis of two-handed object interactions from text as well as generalization to unseen objects.
@@ 2.2. Diffusion for Motion Synthesis
Diffusion models (Sohl-Dickstein et al., 2015) have gained popularity in many domains, such as image generation (Ho et al., 2020), video generation (Yang et al., 2023), audio synthesis (Kong et al., 2021) or hand reconstruction (Ye et al., 2023b, a). More recently, diffusion models have also been adapted to human motion synthesis (Tevet et al., 2023; Zhang et al., 2022b). Various improvements have been proposed, such as integrating physics (Yuan et al., 2023), scene-awareness (Huang et al., 2023), increasing efficiency by diffusing in a pre-trained latent space (Chen et al., 2023a), or the composition of multiple actions (Athanasiou et al., 2022). To enable more controllability during motion synthesis, a two-stage framework, GMD, that guides the diffusion model towards target objectives at inference time is proposed in Karunratanakul et al. (2023). However, GMD contains design choices specific to human motion generation, by first generating a 2D root trajectory over the whole sequence, and then generating corresponding full-body poses, which is not directly transferable to HOI. While these works focus on human motion in isolation from objects, InterDiff (Xu et al., 2023) generates human-object interactions via diffusion models. Their work is improved upon by recent concurrent works (Diller and Dai, 2023; Li et al., 2023; Peng et al., 2023) that leverage contact-based predictions in combination with inference time-guidance to improve the interaction quality. However, these works focus on full-body motions and neglect intricate hand-object interactions. In contrast, we focus on the detailed interactions of fingers and objects. Most similar to ours, the concurrent work MACS (Shimada et al., 2023) proposes a diffusion model for hand-object motion synthesis. Their focus is on synthesizing interactions for a single object with varying mass, while our focus is and generalization to unseen shapes and interaction synthesis conditioned on text.
3. Denoising Diffusion Probabilistic Models
Denoising Diffusion Probabilistic Models (DDPM) (Sohl-Dickstein et al., 2015; Ho et al., 2020) aim to model the probability distribution of a given dataset and can model complex distributions of images (Rombach et al., 2022), videos (Ho et al., 2022) or time-series data such as motion sequences (Tevet et al., 2023).
The diffusion model involves a forward process, which is a Markov chain and consists of sequentially adding Gaussian noise to the data . Each step in the forward noising process results in a distribution, , where is a hyper-parameter and , where is the final number of steps. Alternatively, can be expressed in terms of as
where, .
The reverse process, also called the denoising process is another Markov chain with learned Gaussian transitions starting at . The denoised distribution at step is given by, , where are the model parameters to be learnt, is the mean, and is the variance. The process of data generation involves sequentially drawing samples from this distribution. The variance is usually set to and which is implemented using an auto-encoder trained to match the mean of forward process posterior which is another Gaussian represented as, , where
and .
The auto-encoder could directly estimate the mean or in Eq. 2, or the noise which is used in the noising process of Eq. 1. When estimating the noise, the denoised mean is obtained as , where is the auto-encoder output. The auto-encoder is trained to minimize
where, is either , , or .

### v2

# DiffH2O: Diffusion-Based Synthesis of Hand-Object Interactions from Textual Descriptions

@@ Abstract.
We introduce DiffH2O, a new diffusion-based framework for synthesizing realistic, dexterous hand-object interactions from natural language. Our model employs a temporal two-stage diffusion process, dividing hand-object motion generation into grasping and interaction stages to enhance generalization to various object shapes and textual prompts. To improve generalization to unseen objects and increase output controllability, we propose grasp guidance, which directs the diffusion model towards a target grasp, seamlessly connecting the grasping and interaction stages through a motion imputation mechanism. We demonstrate the practical value of grasp guidance using hand poses extracted from images or grasp synthesis methods. Additionally, we provide detailed textual descriptions for the GRAB dataset, enabling fine-grained text-based control of the model output. Our quantitative and qualitative evaluations show that DiffH2O generates realistic hand-object motions from natural language, generalizes to unseen objects, and significantly outperforms existing methods on a standard benchmark and in perceptual studies.
@@ 1. Introduction
Modeling and understanding hand-object interactions (HOI) is an important and crucial task to empower machines to interact with and assist humans. Being able to seamlessly generate hand-object motions holds the promise to enable synthetic data generation (Leonardi et al., 2023), assist robots in training human-robot interactions in simulation, or enhance realism in virtual manipulation experiences. For example, in virtual reality (VR), hand interactions still often rely on heuristics and controllers that simply attach objects to the hand according to pre-defined grasps. Being able to faithfully produce object manipulation based on an input signal such as text or a few past frames of hand and object poses could largely increase the immersiveness of such interactions.
@@
Generating realistic hand-object interactions in 3D is challenging as the resulting motions are required to be plausible in several different aspects. First, the motions must be plausible in terms of geometry, such that hand and object intersections are minimized and the grasp appears stable. Second, the motions must be plausible in terms of semantics, such as the hands respect natural object affordances (e.g., a cup is grasped by its handle and not flipped upside down). Third, the motions must be plausible in terms of time, such as the hand and object motions are synchronized and the dynamics appear natural. Another challenge in generating HOI comes from the limited scale of existing hand-object datasets, which are around 10x smaller than human motion datasets (Mahmood et al., 2019) and 1000x smaller than image datasets (Deng et al., 2009).
@@
Several recent works have successfully leveraged diffusion models for full-body human motion generation but either neglect objects (Tevet et al., 2023; Karunratanakul et al., 2023) or focus only on coarse motion of larger objects (e.g., a chair) without finger predictions (Xu et al., 2023a). Another method, IMoS (Ghosh et al., 2023), generates HOI with smaller objects (e.g., a cup or stapler) by first predicting upper-body motions and then optimizing object poses in post-processing. However, IMoS assumes that the object is in-hand from the beginning, handles only objects seen during training, and, as shown in our experiments, often yields artifacts (e.g., hand-object interpenetration or non-plausible contacs). In this paper, we aim to learn a model that generates natural, fine-grained hand-object interactions and generalize to objects unseen during training.
@@
Towards this goal, we contribute a novel diffusion-based framework, DiffH2O, that generates hand-object interactions based on text prompts and geometry of the object. We propose several techniques to deal with the challenges of data scarcity and object generalization. Decoupling Grasping and Interaction: Hand-object interactions can be invariably split into a grasping phase where the hand(s) approaches the static object for grasping, and an interaction phase where the object is manipulated based on an action intent. For example, drinking from a cup can be performed with multiple grasps, and given a grasp, multiple actions are possible. We use this simple observation to obtain a grasping and an interaction diffusion model, and introduce an inpainting technique called subsequence imputing to allow continuity between the two outputs. In contrast to previous diffusion frameworks that employ coarse-to-fine (trajectory-to-full motion) motion generation (Karunratanakul et al., 2023), we split the task temporally and propose a temporal two-stage diffusion model. Pose Representation: We introduce a compact representation that models hands in parametric space. To couple the hands to the object, we represent the global hand positions relative to the normalized object position at the initial frame, and further include distance information between the 3D hand joints and the object surface. While encoding distances helps reason about local shape, representing coordinates relative to the initial frame allow us to better adjust to unseen object geometry compared to per-frame object-relative pose representation used in existing work (Christen et al., 2022). Controllability: We provide two ways to add more control to our diffusion models’ output. First, we propose control through fine-grained textual descriptions. To this end, we contribute textual descriptions to the GRAB dataset (Taheri et al., 2020) and show that they enable more robustness to unseen test prompts and increase controllability, e.g., by allowing to prompt the interacting hand and the action to perform. Second, when a target grasp reference pose for a given object is available, we propose to leverage it to connect the grasping and interaction phases by guiding our two diffusion models at inference time. The grasp reference provides a prior about how and with which hand an object should be grasped. It can either be obtained from from an image-based pose estimator (Pavlakos et al., 2024) or grasp synthesis (Zhang et al., 2024a) as we show in our experiments in Section 5.7.
@@
To demonstrate the effectiveness of our approach, we run perceptual studies and experiments on the GRAB dataset (Taheri et al., 2020). We first compare against IMoS (Ghosh et al., 2023). Our results indicate that our method generates better results across a variety of physics-based and motion metrics. To justify our technical contributions, we compare our final model against human-body diffusion baselines by adapting these methods to HOI. We show that we generate motions of higher quality while enabling more control of the outputs reaching feasible poses. Furthermore, we demonstrate that detailed textual descriptions increase robustness to unseen texts and offer better control of the diffusion model than training with heuristics-generated text descriptions. Decoupling our method into two stages and grasp guidance brings better generalization and high practical value to DiffH2O through controllability, e.g., for downstream tasks such as animation or generating synthetic data at scale. We show this with two applications for grasp guidance: i) using an image-based pose estimator (Pavlakos et al., 2024) and ii) leveraging pre-generated grasp poses (Zhang et al., 2024a) to define target reference grasps, and consequently sampling multiple different actions through text in the interaction stage.
We contribute the following: (i) To the best of our knowledge, DiffH2O is the first method that synthesizes hand-object interactions on unseen objects from textual descriptions. (ii) A temporal two-stage diffusion process that splits generation into grasping and action-based interaction, which improves generalization to different textual prompts. (iii) Grasp guidance and subsequence imputing that can be applied at inference to the diffusion process to increase the controllability of outputs and improve generalization to unseen objects. (iv) Detailed textual descriptions to the GRAB dataset. Code and data will be made public upon acceptance***For code and data, see our project website here: https://diffh2o.github.io.
@@ 2.1. Hand-Object Synthesis
Recent efforts in interaction synthesis have largely been driven by the surge of high-quality human-object interaction datasets (Taheri et al., 2020; Brahmbhatt et al., 2019; Hampali et al., 2020; Brahmbhatt et al., 2020; Chao et al., 2021; Kwon et al., 2021; Fan et al., 2023; Liu et al., 2022; Hasson et al., 2019; Yang et al., 2022; Zhan et al., 2024; Wang et al., 2023). Some studies focus on generating coarse full-body object interactions (Zhang et al., 2022a; Wang et al., 2021; Lee and Joo, 2023; Hassan et al., 2023; Luo et al., 2021), such as carrying boxes. FLEX (Tendulkar et al., 2023) trains a hand and body pose prior and optimizes the priors to achieve diverse, static full-body grasps. GOAL (Taheri et al., 2022) and SAGA (Wu et al., 2022) use CVAEs to generate approaching motions for full-body grasps, whereas TOHO (Li et al., 2024b) models both approaching and manipulation tasks via neural implicit representations. Similarly, Braun et al. (2024) model the full range of motion, leveraging physics simulation and reinforcement learning (RL). In contrast, we focus on fine-grained HOI and generalization to unseen objects. Among methods that model full-body object interaction, closest to our work is IMoS (Ghosh et al., 2023), a two-stage method to generate HOI on seen objects based on language prompts. Starting from a grasping state, they first generate body motions and then optimize for object trajectories using a heuristics-based optimization to model HOI. By contrast, our model directly predicts the hand and object poses, models both approaching and manipulation, and generalizes to unseen objects.
@@
Another line of research focuses on generation of hand-object interaction sequences in isolation from full-body due to wide applications in VR and the need for a dedicated model for fine-grained details of finger motion. One prominent solution is to turn to physical simulation and RL, e.g., by learning dexterous manipulation tasks in simulation from full human demonstration data either collected via teleoperation (Rajeswaran et al., 2018) or from videos (Garcia-Hernando et al., 2020; Qin et al., 2021). Mandikal and Grauman (2021) propose a reward function that incentivizes dexterous robotic hand policies to grasp in the affordance region of objects. Li et al. (2024a) propose a solution for grasping multiple objects at once. She et al. (2022) propose a new dynamic state representation to generate dexterous grasps, whereas other works leverage implicit representation (Karunratanakul et al., 2020) or force-closure (Liu et al., 2021). D-Grasp (Christen et al., 2022) and Unidexgrasp (Xu et al., 2023b) leverage RL and physics simulation to generate diverse hand-object interactions from sparse reference inputs. ArtiGrasp (Zhang et al., 2024b) extends this to two-handed grasps and generates articulated object motions. Unidexgrasp++(Wan et al., 2023) proposes a solution that transfers to vision-based inputs, which may be used for robotic grasping. These works focus on achieving stable grasping, whereas we model both the grasping and manipulation of objects.
@@
To model HOI, another solution is to rely on purely data-driven frameworks. For example, (Zhou et al., 2022; Liu and Yi, 2024; Taheri et al., 2024) propose methods that enable denoising hand-poses from noisy sequences of hand-object poses. Ye and Liu (2012) predict the local hand pose given full body and object motion. Similarly, ManipNet (Zhang et al., 2021) predicts local hand poses for two-handed interactions based on wrist-object trajectories. Given object motion of articulated objects, CAMS (Zheng et al., 2023) predicts one-handed poses that align with the object motion, whereas Chen et al. (2023b) focus on grasping objects using dexterity in the environment. All these works either assume a hand-object trajectory, focus only on one hand, ignore semantic action information, or do not generalize well to unseen objects. In contrast, we propose a framework that allows the synthesis of two-handed object interactions from text as well as generalization to unseen objects.
@@ 2.2. Diffusion for Motion Synthesis
Diffusion models (Sohl-Dickstein et al., 2015) have gained popularity in many domains such as image generation (Ho et al., 2020), video generation (Yang et al., 2023), audio synthesis (Kong et al., 2021) or hand reconstruction (Ye et al., 2023b, a). They have also been adapted to human motion synthesis (Tevet et al., 2023; Zhang et al., 2022b). Various improvements have been proposed by integrating physics (Yuan et al., 2023), scene-awareness (Huang et al., 2023), increasing efficiency by diffusing in a pre-trained latent space (Chen et al., 2023a), or the composition of multiple actions (Athanasiou et al., 2022). To enable more controllability during motion synthesis, Karunratanakul et al. (2023) propose GMD that guides the diffusion model towards target objectives at inference time. However, GMD contains design choices specific to human motion generation, by first generating a 2D root trajectory over the whole sequence, and then generating corresponding full-body poses, which is not directly transferable to HOI. While these works focus on human motion in isolation from objects, InterDiff (Xu et al., 2023a) generates human-object interactions via diffusion models. Their work is improved upon by recent concurrent works (Diller and Dai, 2023; Li et al., 2023; Peng et al., 2023) that leverage contact-based predictions in combination with inference time-guidance to improve the interaction quality. However, these works focus on full-body motions and neglect intricate hand-object interactions. In contrast, we focus on the detailed interactions of fingers and objects. Most similar to ours, the concurrent work MACS (Shimada et al., 2023) proposes a diffusion model for hand-object motion synthesis. Their focus is on synthesizing interactions for a single object with varying mass, while our focus is on generalization to unseen shapes and interaction synthesis conditioned on text.
3. Preliminary: Diffusion Models
Denoising Diffusion Probabilistic Models (DDPM) (Sohl-Dickstein et al., 2015) model the probability distribution of a given dataset for images (Rombach et al., 2022), videos (Ho et al., 2022) or motion sequences (Tevet et al., 2023). The diffusion model involves a forward process, which is a Markov chain consisting of sequentially adding Gaussian noise to the data, and a reverse process to denoise the data gradually. The forward process results in a distribution, , where a clean input is gradually noised to , with as the total number of diffusion steps. The reverse process models the probability distribution, , where the noisy input is denoised into , which is modeled via a neural network with parameters . The process of data generation involves sequential denoising and noising steps. The denoising is modeled in auto-encoder fashion predicting either directly the denoised output, ; the mean, , or the noise, , used in the noising process. The auto-encoder is trained to minimize